OpenStudy (is3535):

WILL FAN & MEDAL HELP What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {14, 18, 21, 15, 17} Data Set 2: {15, 17, 22, 20, 16}

OpenStudy (is3535):

@Kash_TheSmartGuy

OpenStudy (is3535):

@Michele_Laino

OpenStudy (is3535):

@phi

OpenStudy (is3535):

@surjithayer

OpenStudy (phi):

did you find the mean of both data sets yet ? (the average)

OpenStudy (is3535):

17 and 18

OpenStudy (is3535):

@phi

OpenStudy (anonymous):

do you know how to find mad?

OpenStudy (phi):

for the mad, on data set 1, first find the median. that means sort the numbers, and find the "middle" number on the sorted list.

OpenStudy (is3535):

i got 17 and 18 for the mean

OpenStudy (phi):

Let's just do data set 1. the mean is 17 I think the median is also 17: 14, 18, 21, 15, 17 becomes 14, 15, 17, 18, 21

OpenStudy (phi):

notice 17 is the middle number after you sort the numbers small to big. to find the MAD, find the difference between the median and each number on the list. Always do bigger number minus smaller number (in other words, we want the absolute value) what new list of numbers do you get ?

OpenStudy (is3535):

it the same

OpenStudy (phi):

yes, the median is also 17. now subtract 17 from each number on the list and take the absolute value. what is the new list?

OpenStudy (is3535):

1. is 3 14 - 17 is 3

OpenStudy (is3535):

like that @phi

OpenStudy (is3535):

or no

OpenStudy (phi):

yes, but we want all five numbers

OpenStudy (is3535):

18-17=1

OpenStudy (is3535):

21-17=4

OpenStudy (is3535):

17-15=2

OpenStudy (is3535):

17-17=0

OpenStudy (is3535):

im done

OpenStudy (phi):

now find the average of those 5 numbers:

OpenStudy (is3535):

all the number are 3 1 4 2 and 0

OpenStudy (phi):

yes, find the average

OpenStudy (is3535):

3 + 1 +4 +2 +0 = 10 divde by 5 = 2

OpenStudy (phi):

yes, now we write the ratio of mean 17 to MAD 2 17/2 or 8.5 that is the "mean to MAD ratio" for data set1

OpenStudy (is3535):

now what

OpenStudy (phi):

now we do data set2 the mean of data set 2 is 18 I think we can use the mean to find the MAD: subtract 18 from each number in data set2 and take the absolute value then find the average of those numbers to get the MAD

OpenStudy (is3535):

18-15=3

OpenStudy (is3535):

18-17=1

OpenStudy (is3535):

22=18=4

OpenStudy (is3535):

20-18=2

OpenStudy (is3535):

18-16=2

OpenStudy (is3535):

im done @phi

OpenStudy (phi):

ok, now what is the average of those?

OpenStudy (is3535):

2.4

OpenStudy (phi):

so the MAD is 2.4 the mean-to-MAD ratio is 18/2.4

OpenStudy (is3535):

7.5

OpenStudy (phi):

I think those are the answers: 8.5 for data set 1 and 7.5 for data set 2

OpenStudy (is3535):

i can only choose one answer

OpenStudy (is3535):

@phi

OpenStudy (is3535):

i can only choose one answer

OpenStudy (phi):

oh. what are the choices?

OpenStudy (is3535):

i have no choices

OpenStudy (is3535):

now what @phi

OpenStudy (phi):

I am assuming that each data set has its own mean-to-MAD ratio, so there are two answers (one for each data set) If there is only one answer, then I don't understand the problem. If your book or notes have an example of this type of problem, maybe I can figure it out.

OpenStudy (is3535):

give me a sec

OpenStudy (phi):

my only other thought is to lump all the numbers together and find the average of all 10 numbers, and the MAD, and then the mean-to-MAD ratio.

OpenStudy (is3535):

it its 7.5 two data sets, expressed as a decimal to the nearest tenth? two data sets